# Fair Resource Allocation in Weakly Coupled Markov Decision Processes

### v1

@@ Appendix B Proofs of Section 3
We start this section with some preliminary results regarding 1) the effect of replacing a policy with one with permuted indices on the value function of a symmetric WCMDP (Section B.1); and 2) a well known result due to Puterman, (2005) on the equivalency between stationary policies and occupancy measures (Section B.2). This is followed with the proof of Lemma 3.3 and our main Theorem 3.4 in sections B.3 and B.4 respectively.
@@ B.1 Value Function under Permuted Policy for Symmetric WCMDP
where we used the fact that the sub-MDPs are identical so that $p^{(N)}(\bm{s}^{\prime}|\bm{s},\bm{a})=p^{(N)}(Q\bm{s}^{\prime}|Q\bm{s},Q\bm{a})$.
@@ B.2 Mapping between stationary policies and occupancy measures
By replacing the term in parentheses as the occupancy measure $q_{\pi}(\bm{s},\bm{a})$ in equation 5:
$\bm{V}^{\pi}_{0}=\sum_{\bm{s}\in\mathcal{S}^{(N)}}\sum_{\bm{a}\in\mathcal{A}^{(N)}}q_{\pi}(\bm{s},\bm{a})\bm{r}(\bm{s},\bm{a}).$
This completes the proof. $\square$
@@
From Lemma B.2, a stationary policy $\bar{\pi}$ can be constructed such that its occupancy measure matches $\bar{q}(\bm{s},\bm{a})$. Namely,
@@
Further, given any ${\bm{w}}$, let us denote with $\pi_{{\bm{w}}}^{*}$ any optimal policy to the GGF problem with ${\bm{w}}$ weights. One can establish that:
@@ Appendix C Count Aggregation MDP
Reward Function  The average reward for all users is defined as
@@
Transition Probability  The transition probability $p^{(N)}_{\phi}(\bm{x}^{\prime}|\bm{x},\bm{u})$ is the probability that the number of MDP in each state passes from $\bm{x}$ to $\bm{x}^{\prime}$ given the action counts $\bm{u}$. We define the pre-image $f^{-1}(\bm{x}^{\prime})$ as the set containing all elements $\bm{s}^{\prime}\in\mathcal{S}^{(N)}$ that maps to $\bm{x}^{\prime}$, then $p^{(N)}_{\phi}(\bm{x}^{\prime}|\bm{x},\bm{u})=\sum_{\bm{s}^{\prime}\in f^{-1}(\bm{x}^{\prime})}p^{(N)}(\bm{s}^{\prime}|f^{-1}(\bm{x}),g_{s}^{-1}(\bm{u}))$.
Given the equivalence of transitions within the pre-image set, for arbitrary state-action pair $(\bm{s},\bm{a})\in\phi^{-1}(\bm{x},\bm{u})$, the probability of transitioning from $\bm{x}$ to $\bm{x}^{\prime}$ under action $\bm{u}$ is the sum of the probabilities of all the individual transitions in the original space that correspond to this count aggregation transition. By using the transition probability in the product space, we obtain
@@ D.1 Optimal Solutions to the GGF-WCMDP problem
The dual linear programming formulation for addressing the multi-objective joint MDP can be naturally extended to the context of vector optimization:
@@ D.2 Solving Count Aggregation MDP by the Dual LP Model
Since the exact model for the count aggregation MDP $\mathcal{M}_{\phi}$ is obtained (Appendix C), a dual linear programming model is formulated following section 6.9.1 of Puterman, (2005), but with count aggregation representation to solve (4):
$\begin{array}[]{rl}\max&\sum\limits_{\bm{x}\in\mathcal{S}^{(N)}_{f}}\sum\limits_{\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}}\bar{r}_{\phi}(\bm{x},\bm{u})q_{\phi}(\bm{x},\bm{u})\\ \mathrm{s.t.}&\sum\limits_{\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}}q_{\phi}(\bm{x},\bm{u})-\gamma\sum\limits_{\bm{x}^{\prime}\in\mathcal{S}^{(N)}_{f}}\sum\limits_{\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}}q_{\phi}(\bm{x}^{\prime},\bm{u})p^{(N)}_{\phi}(\bm{x}|\bm{x}^{\prime},\bm{u})=\mu^{(N)}_{f}(\bm{x}),\,\forall\bm{x}\in\mathcal{S}^{(N)}_{f}\\ &q_{\phi}(\bm{x},\bm{u})\geq 0\hfill\forall\bm{x}\in\mathcal{S}^{(N)}_{f},\forall\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}\end{array}$ (18)
@@ Appendix E Model simulator
In the learning setting, the deep RL agent interacts with a simulated environment through a model simulator (Algorithm 2), which leads to the next states $\bm{x}^{\prime}$ and the average reward $\bar{r}_{\phi}$ across all coupled MDPs.
@@ F.1 Parameter setting
This section details the construction of the components used to generate the test instances based on Akbarzadeh and Mahajan, (2019), including cost function, transition matrix, and reset probability. This experiment uses a synthetic data generator implemented on our own that considers a system with $S$ states and binary actions ($A=2$). There are two possible actions: operate (passive action 0) or replace (active action 1). After generating the cost matrix of the size $\mathcal{S}^{(N)}\cdot\mathcal{A}^{(N)}\cdot N$, we normalize the costs to the range [0, 1] by dividing each entry by the maximum cost over all state action pairs. This ensures that the discounted return always falls within the range $[0,\frac{1}{1-\gamma}]$.
Cost function   The cost function $c(s)$ for $s\in[S]$ can be defined in five ways, each representing a different cost structure associated with each state:
  • Linear: In this case, the cost function $c(s)=s-1$ increases linearly with the state, which means higher states (worse conditions) incur higher costs.
  • Quadratic: The cost function $c(s)=(s-1)^{2}$ applies a more severe penalty for higher states compared to linear cost function.
  • Exponential: The cost function $c(s)=e^{s-1}$ imposes exponentially increasing costs.
  • Ratio of the Replacement Cost Constant (RCCC): The cost function $1.5(S-1)^{2}$ is constant and based on a ratio of 1.5 to the maximum quadratic cost.
  • Random: The cost function $c(s)$ is generated randomly within a specified range of [0,1].
Transition function   The transition matrix for the deterioration action is constructed as follows. Once the machine reaches the $S$-th state, it remains in that state indefinitely until being reset by an active action with a probability of 1. For the $s$-th state $s\in[S-1]$, the probability of remaining in the same state at the next step is given by a model parameter $p_{m}\in[0,1]$, and the probability to the $s+1$-th state is denoted by $1-p_{m}$.
@@ F.2 Chosen Cost Structures
    Quadratic-RCCC: In contrast to scenario ii), operational costs increase quadratically with machine age, while replacement costs remain constant and always higher than operational costs. This setup is typical for high-valued machinery, where the cost of replacement can be significant compared to operational expenses.
F.3 Hyperparameters
In our experimental setup, we chose PPO algorithm to implement the count-proportion based architecture. The hidden layers are fully connected and the Tanh activation function is used. There are two layers, with each layer consisting of 64 units. The learning rates for the actor is set to $5\times 10^{-4}$ and the critic is set to $3\times 10^{-4}$.
F.4 Additional Results on LP Solving Times
The results in Tables 3 and 4 provide details on solving the GGF-LP model and the Count dual LP model on the Quadratic-RCCC instances as the number of machines $N$ increases from 2 to 7. The state size is set to $S=3$, the action size to $A=2$, and the resource to $b=1$. The first block of the tables shows the number of constraints and variables. The second block provides the model solving times, with the standard deviations listed in parentheses. The results are based on 5 runs. Notice that the LP solve time inludes pre-solve, the wallclock time, and post-solve times. The wallclock time is listed separately to highlight the difference from the pure LP Solve time, but not included in the total time calculation.
Table 3: Statistics for GGF-LP Model (3).
@@ F.4 Additional Results on LP Solving Times
Table 4: Statistics for Count dual LP Model (D.2).

### v2

@@ Appendix B PROOFS OF SECTION 3
We start this section with some preliminary results regarding 1) the effect of replacing a policy with one that has permuted indices on the value function of a symmetric WCMDP (Section B.1); and 2) a well-known result from Puterman, (2005) on the equivalency between stationary policies and occupancy measures (Section B.2). This is followed by the proof of Lemma 3.3, which helps establish our main result, Theorem 3.4 in Section B.4.
@@ B.1 Value Function under Permuted Policy for Symmetric WCMDP
where we use the fact that the sub-MDPs are identical so that $p^{(N)}(\bm{s}^{\prime}|\bm{s},\bm{a})=p^{(N)}(Q\bm{s}^{\prime}|Q\bm{s},Q\bm{a})$.
@@ B.2 Mapping Between Stationary Policies and Occupancy Measures
Replacing the term in parentheses as the occupancy measure $q_{\pi}(\bm{s},\bm{a})$ in (5) leads directly to equation 7, which completes the proof. $\square$
@@
From Lemma B.2, a stationary policy $\bar{\pi}$ can be constructed such that its occupancy measure matches $\bar{q}(\bm{s},\bm{a})$:
@@
Furthermore, given any ${\bm{w}}$, let us denote with $\pi_{{\bm{w}}}^{*}$ any optimal policy to the GGF problem with ${\bm{w}}$ weights. One can establish that:
@@
B.5 Extension to Other Fairness Measures
The utilitarian reduction (Theorem 3.4) can be extended to a broader scope of fairness measures, where we replace the GGF measure in optimization problem (2) and define the $\rho$-WCMDP problem accordingly.
Corollary B.4.1.
Let $\rho:\mathcal{V}\rightarrow\mathbb{R}$, with $\mathcal{V}\subseteq\mathbb{R}^{N}$, be a fairness measure that satisfies:
  • (Concavity): The set $\mathcal{V}$ is convex and $\forall{\bm{v}},{\bm{w}}\in\mathcal{V}$, and $\theta\in[0,1]$, $\rho[\theta{\bm{v}}+(1-\theta){\bm{w}}]\geq\theta\rho[{\bm{v}}]+(1-\theta)\rho[{\bm{w}}]$
  • (Permutation invariance): $\forall{\bm{v}}\in\mathcal{V}$ and all $Q\in\mathcal{G}^{N}$, both $Q{\bm{v}}\in\mathcal{V}$ and $\rho[{\bm{v}}]=\rho[Q{\bm{v}}]$
  • (Constant vector invariant) $\forall\bar{v}\in\mathbb{R}\cap\mathcal{V}$, $\rho[\bar{v}{\bm{1}}]=\bar{v}$.
For a symmetric WCMDP such that $\bm{V}_{0}^{\pi}\in\mathcal{V}$ for all $\pi$, let $\Pi^{*}_{\mbox{\scriptsize U, PI}}$ be the set of optimal policies for the utilitarian approach that is permutation invariant, then $\Pi^{*}_{\mbox{\scriptsize U, PI}}$ is necessarily non-empty and all $\pi^{*}_{\mbox{\scriptsize U, PI}}\in\Pi^{*}_{\mbox{\scriptsize U, PI}}$ satisfies
$\rho[\bm{V}_{0}^{\pi_{\mbox{\tiny{U, PI}}}^{*}}]=\max_{\pi}\rho\left[\bm{V}^{\pi}_{0}\right].$
Proof. Similar to the proof of Theorem 3.4, we define the utilitarian fairness measure as
$\rho_{U}[{\bm{v}}]=\frac{1}{N}\sum_{n=1}^{N}v_{n}.$
Defining the optimal policy to the $\rho_{U}$-WCMDP problem with utilitarian objective as $\pi^{*}_{U}\in\arg\max\limits_{\pi}\rho_{U}[\bm{V}^{\pi}_{0}]$ and with Lemma 3.3, we can construct a permutation invariant policy $\bar{\pi}^{*}_{U}$ satisfying
$\bar{V}_{0}^{\pi^{*}_{U}}{\bm{1}}=\bm{V}_{0}^{\bar{\pi}^{*}_{U}}\in\mathcal{V},$ (13)
with $\bar{V}_{0}^{\pi^{*}_{U}}:=\frac{1}{N}\sum_{n=1}^{N}V_{0,n}^{\pi^{*}_{U}}$. Then we have that
$\rho_{U}\left[\bm{V}_{0}^{\pi^{*}_{U}}\right]=\bar{V}_{0}^{\pi^{*}_{U}}=\rho\left[\bar{V}_{0}^{\pi^{*}_{U}}\right]=\rho\left[\frac{1}{N}\sum_{n=1}^{N}V_{0,n}^{\pi^{*}_{U}}{\bm{1}}\right]=\rho\left[\bm{V}_{0}^{\bar{\pi}^{*}_{U}}\right],$
where we exploit $\rho[\bar{v}{\bm{1}}]=\bar{v}$ for all $\bar{v}\in\mathbb{R}\cap\mathcal{V}$. Furthermore, let $\pi^{*}$ be any optimal policy to the $\rho$-WCMDP problem. One can establish that:
$\rho[\bm{V}_{0}^{{\pi}^{*}}]\geq\rho\left[\bm{V}_{0}^{\bar{\pi}^{*}_{U}}\right]=\rho_{U}\left[\bm{V}_{0}^{\pi^{*}_{U}}\right]\geq\rho_{U}\left[\bm{V}_{0}^{{\pi}^{*}}\right].$ (14)
By Jensen’s inequality and the fact that $\rho[\cdot]$ is concave, it holds that $\rho_{U}[\bm{V}_{0}^{{\pi}^{*}}]\geq\rho[\bm{V}_{0}^{{\pi}^{*}}]$ since for all ${\bm{v}}\in\mathcal{V}$, we have that
$\rho[{\bm{v}}]=\frac{1}{N!}\sum_{Q\in\mathcal{G}^{N}}\rho[Q{\bm{v}}]\leq\rho[\frac{1}{N!}\sum_{Q\in\mathcal{G}^{N}}Q{\bm{v}}]=\rho[\frac{1}{N}\sum_{n=1}^{N}v_{n}{\bm{1}}]=\frac{1}{N}\sum_{n=1}^{N}v_{n}=\rho_{U}[{\bm{v}}],$
where we use permutation invariance of $\rho$, followed with its concavity and its contant vector invariance. The inequalities in (14) should therefore all reach equality:
$\rho[\bm{V}_{0}^{\pi^{*}}]=\rho\left[\bm{V}_{0}^{\bar{\pi}^{*}_{U}}\right]=\rho_{U}\left[\bm{V}_{0}^{\pi^{*}_{U}}\right].$
The rest of the argument follows exactly as in the proof of Theorem 3.4 (see Appendix B.4). $\square$
Now, we comment that the expected utility model $\rho[{\bm{v}}]=u^{-1}\left({\frac{1}{N}}\sum_{n=1}^{N}u(v_{n})\right)$, where $u(\cdot)$ is a monotone and concave function, satisfies the three properties defined in Corollary B.4.1 and is a natural framework to measure fairness in resource allocation problems as discussed in Bertsimas et al., (2012). The concavity of $u(\cdot)$ reflects a decreasing marginal utility on the allocated resource to an individual. This property promotes equitable distributions of resources by discouraging disparities in utility. A notable instance of this model is $\alpha$-fairness (Mo and Walrand,, 2000; Ju et al.,, 2023), which is parameterized by $\alpha>0$ and takes the form
$u_{\alpha}(v):=\left\{\begin{array}[]{cl}\log(v)&\mbox{if $\alpha=1$,}\\ \frac{v^{1-\alpha}}{1-\alpha}&\mbox{if $\alpha\neq 1$.}\end{array}\right.$
The domain of $\rho$ is restricted to non-negative if $\alpha\neq 1$ and strictly positive otherwise. This function covers a range of fairness objectives, from the proportional fairness ($\alpha$ = 1) to the max-min fairness ($\alpha\rightarrow\infty$).
@@ Appendix C COUNT AGGREGATION MDP
Reward Function  The average reward for all sub-MDPs is defined as
@@
Transition Probability  The transition probability $p^{(N)}_{\phi}(\bm{x}^{\prime}|\bm{x},\bm{u})$ is the probability that the number of sub-MDPs in each state passes from $\bm{x}$ to $\bm{x}^{\prime}$ given the action counts $\bm{u}$. We define the pre-image $f^{-1}(\bm{x}^{\prime})$ as the set containing all elements $\bm{s}^{\prime}\in\mathcal{S}^{(N)}$ that map to $\bm{x}^{\prime}$, then $p^{(N)}_{\phi}(\bm{x}^{\prime}|\bm{x},\bm{u})=\sum_{\bm{s}^{\prime}\in f^{-1}(\bm{x}^{\prime})}p^{(N)}(\bm{s}^{\prime}|f^{-1}(\bm{x}),g_{s}^{-1}(\bm{u}))$.
Given the equivalence of transitions within the pre-image set, for an arbitrary state-action pair $(\bm{s},\bm{a})\in\phi^{-1}(\bm{x},\bm{u})$, the probability of transitioning from $\bm{x}$ to $\bm{x}^{\prime}$ under action $\bm{u}$ is the sum of the probabilities of all the individual transitions in the original space that correspond to this count aggregation transition. By using the transition probability in the product space, we obtain
@@ D.1 Optimal Solutions to the GGF-WCMDP Problem
The dual LP formulation for addressing the multi-objective joint MDP can be naturally extended to the context of vector optimization:
@@ D.2 Solving Count Aggregation MDP by the Dual LP Model
Since the exact model for the count aggregation MDP $\mathcal{M}_{\phi}$ is obtained (Appendix C), a dual LP model is formulated following Section 6.9.1 of Puterman, (2005), but with count aggregation representation to solve (4):
$\begin{array}[]{rl}\max&\sum\limits_{\bm{x}\in\mathcal{S}^{(N)}_{f}}\sum\limits_{\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}}\bar{r}_{\phi}(\bm{x},\bm{u})q_{\phi}(\bm{x},\bm{u})\\ \mathrm{s.t.}&\sum\limits_{\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}}q_{\phi}(\bm{x},\bm{u})-\gamma\sum\limits_{\bm{x}^{\prime}\in\mathcal{S}^{(N)}_{f}}\sum\limits_{\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}}q_{\phi}(\bm{x}^{\prime},\bm{u})p^{(N)}_{\phi}(\bm{x}|\bm{x}^{\prime},\bm{u})=\mu^{(N)}_{f}(\bm{x}),\,\forall\bm{x}\in\mathcal{S}^{(N)}_{f}\\ &q_{\phi}(\bm{x},\bm{u})\geq 0\hfill\forall\bm{x}\in\mathcal{S}^{(N)}_{f},\forall\bm{u}\in\mathcal{A}^{(N)}_{g_{s}}\end{array}.$ (20)
@@ Appendix E MODEL SIMULATOR
In the learning setting, the deep RL agent interacts with a simulated environment through a model simulator (Algorithm 2), which leads to the next state $\bm{x}^{\prime}$ and the average reward $\bar{r}_{\phi}$ across all coupled MDPs.
@@ F.1 Parameter Setting
This section details the construction of the components used to generate the test instances based on Akbarzadeh and Mahajan, (2019), including the cost function, the transition matrix, and the reset probability. This experiment uses a synthetic data generator implemented on our own that considers a system with $S$ states and binary actions ($A=2$), where the two possible actions are to operate or to replace. After generating the cost matrix of the size $\mathcal{S}^{(N)}\cdot\mathcal{A}^{(N)}\cdot N$, we normalize the costs to the range [0, 1] by dividing each entry by the maximum cost over all state action pairs. This ensures that the discounted return always falls within the range $[0,\frac{1}{1-\gamma}]$.
Cost function   The cost function $c(s)$ for $s\in[S]$ can be defined in five ways: 1) Linear: $c(s)=s-1$, where the cost increases linearly with the state index; 2) Quadratic: $c(s)=(s-1)^{2}$ with a more severe penalty for higher states compared to the linear case; 3) Exponential: $c(s)=e^{s-1}$, which leads to exponentially increasing costs; 4) Replacement Cost Constant Coefficient (RCCC): $c(s)=1.5(S-1)^{2}$, which is based on a constant ratio of 1.5 to the maximum quadratic cost; 5) Random: $c(s)$ is randomly generated within the range [0,1].
Transition function   The transition matrix for the deterioration action is constructed as follows. Once the machine reaches the $S$-th state, it remains in that worst state indefinitely until being successfully reset by a replacement action. For the $s$-th state $s\in[S-1]$, the probability of remaining in the same state at the next step is given by a model parameter $p_{m}\in[0,1]$, and the probability of transitioning to the $(s+1)$-th state is $1-p_{m}$.
@@ F.2 Chosen Cost Structures
    Quadratic-RCCC: In contrast to scenario ii), operational costs increase quadratically with machine age, while replacement costs remain constant and always higher than operational costs. This setup is typical for high-valued machinery, where replacement costs can be significant compared to operational expenses.
F.3 Comparison of CP-DRL Algorithms
As presented in Tables 3 and 4, the PPO algorithm consistently shows high-quality performance with low variance compared to TD3 and SAC. This motivated us to choose PPO to implement the CP-DRL algorithm.
Table 3: GGF Scores (Exponential-RCCC)
$N$=2 $N$=3 $N$=4 $N$=5
PPO 14.11 $\pm$ 0.01 13.89 $\pm$ 0.14 13.59 $\pm$ 0.10 13.28 $\pm$ 0.03
TD3 11.83 $\pm$ 2.74 12.30 $\pm$ 1.14 12.63 $\pm$ 0.23 12.62 $\pm$ 0.12
SAC 14.00 $\pm$ 0.05 13.76 $\pm$ 0.03 13.50 $\pm$ 0.05 13.30 $\pm$ 0.02
Table 4: GGF Scores (Quadratic-RCCC)
$N$=2 $N$=3 $N$=4 $N$=5
PPO 16.14 $\pm$ 0.00 16.05 $\pm$ 0.00 15.94 $\pm$ 0.00 15.86 $\pm$ 0.02
TD3 15.98 $\pm$ 0.03 15.75 $\pm$ 0.12 15.56 $\pm$ 0.24 15.51 $\pm$ 0.34
SAC 16.06 $\pm$ 0.01 15.75 $\pm$ 0.08 15.69 $\pm$ 0.04 15.28 $\pm$ 0.05
F.4 Hyperparameters
In our experimental setup, we chose PPO algorithm to implement the count-proportion based architecture. The hidden layers are fully connected and the Tanh activation function is used. There are two layers, with each layer consisting of 64 units. The learning rate for the actor is set to $5\times 10^{-4}$ and the critic is set to $3\times 10^{-4}$. The Vanilla-DRL baseline uses the same network architecture as PPO, with two hidden layers of 64 neurons each, but with a softmax output layer for its stochastic policy.
F.5 Additional Results on LP Solving Times
The results in Tables 5 and 6 provide details on solving the GGF-LP model and the count dual LP model on the Quadratic-RCCC instances as the number of machines $N$ increases from 2 to 7. The state size is set to $S=3$, the action size to $A=2$, and the resource to $b=1$. The first block of the tables shows the number of constraints and variables. The second block provides the model solving times, with the standard deviations listed in parentheses. The results are based on 5 runs. Notice that the LP solve time inludes pre-solve, the wallclock time, and post-solve times. The wallclock time is listed separately to highlight the difference from the pure LP Solve time, but not included in the total time calculation.
Table 5: Statistics for GGF-LP Model (3)
@@ F.5 Additional Results on LP Solving Times
Table 6: Statistics for Count Dual LP Model (D.2)